Game_Number Winner Loser Duration
1
Jens
Sven 41
2 Sven
Jens
76
3 Sven
Jens
156
4 Sven
Jens
151
5 Sven
Jens
105
6
Jens
Sven 71
7 Sven
Jens
123
8 Sven
Jens
111
9 Sven
Jens
164
10 Sven
Jens
116
11
Jens
Sven 46
12 Sven
Jens
88
13
Jens
Sven 59
14
Jens
Sven 47
15
Jens
Sven 57
16 Sven
Jens
108
17 Sven
Jens
78
18 Sven
Jens
94
19
Jens
Sven 86
20
Jens
Sven 60

21

Jens
Sven 52
22
Jens
Sven 64
23
Jens
Sven 72
24
Jens
Sven 56
25
Jens
Sven 30
26 Sven
Jens
173
27
Jens
Sven 46
28
Jens
Sven 80
29 Sven
Jens
147
30
Jens
Sven 87
31
Jens
Sven 71
32 Sven
Jens
146
33 Sven
Jens
66
34
Jens
Sven 62
35
Jens
Sven 59
36
Sven
Jens
173
37
Jens
Sven 73
38
Jens
Sven 75
39 Sven
Jens
148
40
Jens
Sven 73
41 Sven
Jens
127
42 Sven
Jens
108
43
Jens
Sven 32
44
Jens
Sven 61
45
Jens
Sven 89
46
Jens
Sven 34
47
Jens
Sven 50
48
Jens
Sven 62
49 Sven
Jens
128
50 Sven
Jens
73
51 Sven
Jens
112
52 Sven
Jens
125
53 Sven
Jens
65
54 Sven
Jens
141
55 Sven
Jens
104
56 Sven
 
Jens
123
57
Jens
Sven 86
58
Jens
Sven 49
59 Sven
Jens
144
60 Sven
Jens
103
61
Jens
Sven 81
62 Sven
Jens
65
63
Jens
Sven 77
64
Jens
Sven 64
65 Sven
Jens
162
66 Sven
Jens
155
67 Sven
Jens
67
68 Sven
Jens
91
69
Jens
Sven 67
70 Sven
Jens
137
71
Jens
Sven 82
72
Jens
Sven 86
73 Sven
Jens
102
74 Sven
Jens
101
75 Sven
Jens
164
76
Jens
Sven 31
77 Sven
Jens
83
78 Sven
Jens
173
79 Sven
Jens
125
80
Jens
Sven 49
81 Sven
Jens
125
82 Sven
Jens
115
83
Jens
Sven
79
84
Jens
Sven 64
85 Sven
Jens
177
86 Sven
Jens
124
87
Jens
Sven 83
88 Sven
Jens
162
89 Sven
Jens
104
90 Sven
Jens
135
91 Sven
Jens
179
92
Jens
Sven 88
93
Jens
Sven 72
94 Sven
Jens
66
95 Sven
Jens
119
96 Sven
Jens
177
97
Jens
Sven 73
98 Sven
Jens
137
99 Sven
Jens
99
100 Sven
Jens
159
  1. For the games Sven won, provide an estimate of the mean time it takes Sven to beat his brother. Similarly, for the games Jens won, provide an estimate of the mean time it takes for Jens to beat his brother.
Answer #1

total games won by sven = 56

total time taken by sven in his games won = 6879

mean time it takes Sven to beat his brother = 6879/56 = 122.839280

total games won by sven = 44

total time taken by sven in his games won = 2826

mean time it takes for Jens to beat his brother. = 2826/44 = 64.227273